You have to solutions. A 5% saline solution and a 20% saline solution. You have to mix them together to make 90 mL of 10% saline solution. How much of each solution do you need
Answer by mananth(16946)

percent ---------------- quantity
Saline solution type I 20.00% ---------------- x ml
Saline solution type II 5.00% ------ 90 - x ml
Mixture 10.00% ---------------- 90
Total 90 ml
20.00% x + 5.00% ( 90 - x ) = 10.00% * 90
20 x + 5 ( 90 - x ) = 900
20 x + 450 - 5 x = 900
20 x - 5 x = 900 - -450
15 x = 450
/ 15
x = 30 ml 20.00% Saline solution type I
60 ml 5.00% Saline solution type II

m.ananth@hotmail.ca